OpenStudy (hlilly2413):

Got to a certain point, need help from here. Exposure to high doses of microwaves can cause damage to cells. Estimate how many moles of photons with a wavelength of 12 cm must be absorbed to raise the temp. of your eye by 3 degrees Celsius. Assume mass of eye = 11 g and specific heat capacity is 4.0J/g*C. Give me a sec, I'll post what I've done so far as a comment.

OpenStudy (hlilly2413):

So far I've found q or how many energy is needed to raise the temperature of my 11 g eye by 3 degrees Celsius. The equation used (mcdeltaT) is as follows (11g)(4.0J/g*C)(3.0C) = 132 J Next I've calculated the energy of a photon by using the equation hc/wavelength: (6.626*10^-34J/s)(3.00*10^8 m/s) / (.12m) = 1.9878*10^-24 J per photon The question is asking how many moles of photons must be absorbed, so do I do 132 J/ 1,9878*10^24J ? Then the Joules cancel and I'm left with...photons...and there are 6.022*10^23 photons per 1 mole so do I just multiply the answer by Avogadro's number and call it a day?

OpenStudy (hlilly2413):

Woof. How much energy** How many energy, oh man. Can you tell I've been working for awhile? :)

OpenStudy (anonymous):

Find total energy needed, energy per photon, then find total photons. Divide that photon number by Avogardo and you will have moles of photons needed.

OpenStudy (anonymous):

@hlilly2413

OpenStudy (hlilly2413):

I have found the energy needed and I have found the energy per photon as stated above. This is probably much more simple than I think, but based on the work that I have done already I'm stuck on the next step. How do I find the total number of photons?

OpenStudy (anonymous):

@hlilly2413 Since you know the energy needed and the amount of energy a photon can transfer, you can find how many photons you need to reach the amount of energy needed. For: Pe = energy per photon Er = energy required to raise the eye temp we can then find, P = number of photons needed The formula would be: \[P*Pe=Er\] Then solve for P: \[P=Er/Pe\] Now you know how many photons needed and just need to put that in mole units. Remember, a mole is simply \[6.02214078 * 10^{23}\] things. Could be marbles, candy bars, anything. In our case it is photons. So, now just find how many times the number of photons needed can go into \[6.02214078 * 10^{23}\] photons, and you will have the moles needed.